OpenStudy (anonymous):

differentiate F(x) = xe^-x^2 = xe^-x^2 (-2x) F(x) =-2x^2 e^-x^2 is that the answer

OpenStudy (tkhunny):

This \(\dfrac{d}{dx}xe^{-x^{2}}\)?

OpenStudy (anonymous):

you've got to use the product rule because there are two functions of "x" multiplying and then during evaluation, you will have to use the chain rule in one of the steps

OpenStudy (anonymous):

\[x(e ^{-x^2})(-2x)+(e ^{-x^2})(1)\] answer \[-2x^2(e ^{-x^2})+e ^{-x^2}\]

OpenStudy (tkhunny):

You may wish to rewrite a bit: \(e^{-x^{2}}(1-2x^{2})\)

OpenStudy (anonymous):

where the one coming from?

OpenStudy (tkhunny):

Multiplication? Consult the Distributive Property of Multiplication over Addition.

OpenStudy (fibonaccichick666):

ok so if tkhunny is correct on the notation, it may help to think of it in terms of a substitution: so \[\frac{d}{dx} xe^{x^2}=x(e^u du)\] applying the chain rule: \[=e^u+xe^udu\] \[=e^{x^2}+x*2x*e^{x^2}\] \[=e^{x^2}[1+2x^2]\] do you understand?

OpenStudy (anonymous):

yes thank you to all of you

OpenStudy (tkhunny):

...well, with a few negative signs in there.